A, B and C are partners sharing profits in the ratio of 5:4:1. C is guaranteed that his share in a year will not be less than ₹5,000. Profit for the year ended 31st March, 2021 is ₹40,000. Deficiency in the guaranteed profit of C is to be borne by B. Deficiency to be borne by B is:


1
Expert's answer
2021-10-26T15:12:16-0400

profit share:

A:

"40000\times\frac{5}{10}=20000"

B:

"40000\times\frac{4}{10}=16000"

C:

"40000\times\frac{1}{10}=4000"


Deficiency in C share:

5000-4000=1000


"DeficiencyA=DeficiencyB=1000\times\frac{1}{2}=500"


profit after deficiency:

A=20000-500=19500

B=16000-500=15000

C=4000+1000=5000